A 4.13 kg block slides down a 6.09 m long ramp that has an angle of inclination of 28.2° from rest. At the bottom of the ramp, the block is travelling 2.98 m/s. What is the work done by friction on the block as it slides down the ramp?



1
Expert's answer
2021-06-25T10:44:44-0400
"W_f=mgh-0.5mv^2\\
W_f=(4.13)(9.8)(6.09)\sin{28.2}-0.5(4.13)(2.98)^2\\W_f=98.1\ J"